# CloudFix: Automated Policy Repair for Cloud Access Control Policies Using Large Language Models

### v1

@@ Abstract
Access control policies are vital for securing modern cloud computing, where organizations must manage access to sensitive data across thousands of users in distributed system settings. Cloud administrators typically write and update policies manually, which can be an error-prone and time-consuming process and can potentially lead to security vulnerabilities. Existing approaches based on symbolic analysis have demonstrated success in automated debugging and repairing access control policies; however, their generalizability is limited in the context of cloud-based access control. Conversely, Large Language Models (LLMs) have been utilized for automated program repair; however, their applicability to repairing cloud access control policies remains unexplored. In this work, we introduce CloudFix, the first automated policy repair framework for cloud access control that combines formal methods with LLMs. Given an access control policy and a specification of allowed and denied access requests, CloudFix employs Formal Methods-based Fault Localization to identify faulty statements in the policy and leverages LLMs to generate potential repairs, which are then verified using SMT solvers. To evaluate CloudFix, we curated a dataset of 282 real-world AWS access control policies extracted from forum posts and augmented them with synthetically generated request sets based on real scenarios. Our experimental results show that CloudFix improves repair accuracy over a Baseline implementation across varying request sizes. Our work is the first to leverage LLMs for policy repair, showcasing the effectiveness of LLMs for access control and enabling efficient and automated repair of cloud access control policies. We make our tool Cloudfix and AWS dataset publicly available.
@@ I Introduction
Cloud compute systems allow administrators to write policies that specify access control rules, which specify how access to private data is governed. Policies are typically written in a convenient language, such as the Amazon Web Services (AWS) Identity and Access Management (IAM) policy language, allowing administrators to write their own policies. However, this is a tedious, complex, and often error-prone process. Moreover, policies must be modified on a regular basis due to the ever changing nature of cloud computing. Creating and maintaining policies is a challenging endeavor and often requires significant technical skills and domain expertise in ensuring access control is correctly implemented. This can potentially lead to smaller organizations or those without the proper resources to forgo best security practices in ensuring access privilege is kept to a minimum, and instead provide broad access in order to ensure their cloud computing services work as intended.
@@
Cloud computing services often provide tools to assist administrators in writing and maintaining policies or monitoring access to resources (e.g., AWs IAM Access Analyzer, AWS IAM Policy Simulator, AWS Cloudtrail service). However, when policies do not function as they are intended, or must be modified to incorporate access to new services or resources, these tools are inadequate when administrators lack the technical expertise to use them. This is additionally compounded by the lack of automated approaches for creating or repairing policies, as existing techniques either do not extend to cloud access control policies or are restricted in the ways they repair policies [Son2013FixMU, peng2017towards, D'Antoni2024, DetecingResolvingPolicyMisconfig, Jayaraman2014AutomatedAA]. What is needed is automated techniques that can reason about the complexities of cloud policies and be able to modify and repair existing policies as permissions and services evolve throughout the cloud ecosystem.
@@
Recently, Large Language Models (LLMs) have seen great success in the field of software engineering [LLM4SESurvey]. They have become invaluable tools for software developers and engineers with their ability to generate code and automate tasks. In particular, LLMs have been shown to be useful for automated program repair, where LLMs generate patches to fix faulty programs [zhang2024systematic, li2025hybrid, hossain2024deep, xia2023automated, campos2025empirical, orvalho2025counterexample]. In the field of access control, LLMs have been shown to be successful in synthesizing policies from natural language descriptions [jayasundara2023sokaccesscontrolpolicy, vatsa2025synthesizingaccesscontrolpolicies, 1269406, EnforcingAccessControlLLMS]. However, their applicability in the context of repairing access control policies has yet to be explored.
@@
In this work, we take the first step towards leveraging Large Language Models for access control policy repair. Given a faulty or misconfigured access control policy and access control requests which must be allowed or denied by the policy, our approach repairs the policy so that it correctly classifies the given set of requests. We implement our approach in a tool called CloudFix and showcase its effectiveness in repairing AWS IAM access control policies. Additionally, we curate the first policy repair dataset for cloud access control policies and make it publicly available for future research endeavors. Specifically, we make the following key contributions:
@@
  • We automate the access control policy repair process using an LLM-driven Fault-localization-guided approach.
@@
  • We experimentally show that our fault-localization techniques enable LLMs to repair faulty policies.
@@
The paper is organized as follows. In Section II we provide the overview access control policies and the policy repair problem. In Section III we introduce CloudFix. In Section IV we describe how we curate a dataset to evaluate CloudFix. In Section V we experimentally validate CloudFix. In Section VI we discuss our findings. In Section VII we discuss the related work. In Section VIII we conclude our paper.
@@ II Overview and Background
In this work we focus primarily on access control policies written in AWS IAM policy language. However, the techniques we present in this work are not specific to the AWS IAM policy language and can be applied to any policy language.
@@ II-A Amazon Web Service Policies
An AWS IAM policy $P$ is a declarative policy consisting of a list of statements describing how access is governed. Policies field access control requests which are either allowed or denied by the policy. An access control request is a tuple $(\delta,a,r,e)$ where $\delta$ is a specific principal performing the access, $a$ is the action to be performed, $r$ is the resource being accessed, and $e$ is the set of environment variables within the request (e.g., IP address, current time, etc). A policy statement $S$ is a 5-tuple (Effect, Principal, Action, Resource, Condition) where:
@@
Initially, all requests are implicitly denied. An access control request is granted access if and only there is at least one statement within the policy which explicitly allows access and no statement in the policy explicitly revokes or denies access. The AWS policy language allows special characters to be used which represent broad access: the ’*’ wildcard character can be used to represent any string, while the ’?’ character can be used to represent any character. Additionally AWS allows identity-based policies which omit the Principal field.
Figure 1: Manual Policy Repair Process. Stage 1 shows an administrator authoring a policy while a user generates access requests. In Stage 2, a cloud evaluation engine validates these requests, correctly allowing one (s3:GetObject) but implicitly denying another (ec2:RunInstances) because no statement permits it. In Stage 3, the administrator manually interprets this feedback, adding a new Allow statement to the policy to fix the issue, and resubmits it for validation.
@@ II-B A Motivating Example
In real-world cloud computing, debugging and repairing access control policies are done collaboratively but largely manually. Administrators write policies in the AWS IAM consoles and attach them to users/roles, or use the IAM Policy Simulator, which still requires manual work. This prompts iterative manual debugging until the desired permissions are allowed and denied. A common yet risky solution for this is to craft overly broad permissions (e.g., action: s3: *) to temporarily prevent requests from being unintentionally denied and attempt to narrow the scope iteratively.
Figure 1 demonstrates how an AWS engineer identifies errors in the policy and attempts to repair it. The process has three states: in Stage 1, the Cloud Security Admin writes an access control policy that gives access to the principal to read any s3 category and object across the account. In the first request, the user is asking to read the object newfile.txt stored in the s3 category my-category by invoking the s3:GetObject action. In the second request, the user is asking to launch new ec2 instances and access the corresponding backend logs in the us-west-2 region by performing the ec2:RunInstances action on the resource arn:aws:ec2:us-west-2:backend/logs. These requests are stored in CloudTrail as access history logs and are accessible to the Admin. Next, the Admin examines the requests through the Cloud Engine to locate the failure points in the policy, as shown in Stage 2. The Cloud Engine validates and returns that the first request is correctly allowed, as the policy has an explicit statement that allows it, and the second request is implicitly denied since there is no statement allowing the request. In Stage 3, the Admin adds an explicit statement in the policy that allows the required request, which was initially denied by the policy. The Admin then uses the Cloud Engine to validate the repaired policy and deploys it.
This manual repair and validation process is tedious, time-consuming, and error-prone. Each time a user’s request is denied or overly allowed, the administrator must inspect the IAM policy line by line, trace the evaluation logic through the Cloud Engine, and cross-check the results against CloudTrail logs to understand why the policy denied or allowed the access request. Manually inspecting the policy to repair it requires redeploying and retesting the policy to confirm that intended requests are still allowed and denied. Moreover, this iterative process heavily relies on human reasoning and repeated trial and error; it becomes increasingly impractical as the number of users, policies, and requests grows.
@@ III CloudFix
In this section we introduce our policy repair framework CloudFix. The framework consists of four steps: (A) Goal Validation, (B) Fault Localization, (C) Prompt Generation, and (D) Repair Synthesizer. Figure 2 demonstrates the overall stages of the method. CloudFix operates over an input access control policy and request specification. The Goal Validator first evaluates the policy against the request specification: if all requests are correctly classified, the policy is considered correct and the framework returns the policy. Otherwise, the process proceeds iteratively. In each iteration, the faulty policy and the requests are fed into the Fault Localizer component that analyzes the failed requests and the current policy to identify statements that are responsible for misclassifications. These faulty statements, together with the policy and requests, are provided to the Prompt Generator, which produces an input prompt for the Repair Synthesizer. The Repair Synthesizer queries an LLM to generate a candidate policy as a potential repair. This candidate policy is validated against the requests by the Goal Validator. If the candidate policy correctly classifies all requests, then the policy has been completely repaired, and the algorithm terminates. Otherwise, the candidate policy becomes the new input to the next iteration of the loop. This process continues until a fully repaired policy is produced, or the iteration threshold is reached. Algorithm 1 presents the complete CloudFix algorithm.
@@ III-A Goal Validator
The Goal Validator determines if the policy correctly classifies all requests in the given specification. Algorithm 2 shows how a request specification, $r$ in a set $R$ is validated against a policy $P$ to test whether $r$ is not allowed by $P$ and determine that $P$ is faulty. To do this, both the policy $P$ and $r$ are encoded into logical formulas suitable for satisfiability checking using the Quacky tool [eiers2022quacky]. The algorithm returns FAIL if at least one request is misclassified. This information is then passed to the Fault Localizer for further identification of statements within $P$ that are responsible for the misclassification of the $r$ in $S$. Encoding the requests and the policies as SMT formulas using an SMT Solver enables us to eliminate the dependency on external policy evaluation engines. We use a maximum of $I$ iterations per policy to reach a repair goal of 100%. This can be defined by the user depending on the available resources. Once the target accuracy or the maximum $I$ is reached, the repair algorithm stops and returns the repaired policy. If the target repair accuracy is not achieved within $I$ iterations, the algorithm retains the repair with the highest accuracy.
@@
4:if $verification=\textsc{pass}$ then
@@
9:for each $(\delta,a,r,e)\in R_{\text{deny}}$ inRR do
@@ III-B Fault Localizer
Once the Goal Validator determines that a policy $\mathit{P}$ is faulty (i.e., it misclassifies requests), then $\mathit{P}$ is passed to the Fault Localizer to identify statements within it that misclassify the requests. We define a request misclassification as allowing a request that is supposed to be denied and/or denying a request that is supposed to be allowed by the policy. We define three types of faults within the policy that are responsible for misclassifications: case 1) faulty statements that explicitly allow requests which should be denied as per the request specification, case 2) statements that explicitly deny requests that should be allowed per the request specification, and case 3) requests which are implicitly denied yet should be allowed. The process is shown in Algorithm 3. Each misclassified request is analyzed with respect to the policy to determine the type of fault that incorrectly classifies it.
Given a policy $\mathit{P}$ and a request set $\mathit{R}$, lines 4 - 17 in Algorithm 3 illustrate how we identify faulty statements within $\mathit{P}$. The $\mathrm{SMTSolver}$ function determines the satisfiability of policy constraints against specific requests using the Z3 solver [z3] for satisfiability checking and the Quacky tool for policy and request encoding. For each misclassified request, to check if it falls into case 1, $\mathrm{SMTSolver}(r,s)$ iteratively checks each allow statement $\mathit{s\in S^{+}}$ to see if it allows request $\mathit{r}$ by creating a minimal policy containing only that statement. Lines 7 and 8 show that if $\mathrm{SMTSolver}(r,s)$ returns $\mathit{SAT}$, we assert that there is an explicit allow statement that is allowing the statement that should have been denied. To check case 2, $\mathrm{SMTSolver}(r,P^{\star}\cup\{s\}))$ combines the universal policy with a deny statement, $\mathit{s}$, to determine if the deny statement explicitly denies $\mathit{r}$. This is shown in lines 11-12. If the solver returns $\mathit{UNSAT}$, it indicates that the deny statement in $\mathit{S}$ is responsible for the incorrect denial of $\mathit{r}$. If this condition fails, case 3 holds, as shown in line 14. Finally, Algorithm 3 returns a list that has a mapping between policy statements and request failures, along with the failure types. This combination of statements, along with the explicit requests they allow, deny, or implicitly deny, is then passed to the Prompt Generator, which is discussed below.
Figure 2: This flowchart details the architecture of the CloudFix framework.
@@ III-C Prompt Generator
Once the Fault Localizer identifies the faulty statements $s$ within $P$, we format the results into a structured natural language prompt for the LLM. The Prompt Generator receives the fault localization report, the list of request specifications $R$, and the policy $P$. It then structures them into a format suitable for LLMs. We annotate the prompts with the iteration number and the accuracy obtained from the Goal Validator to provide contextual feedback to the LLM. In this work, we consider two types of prompts: Base and Fault Localization-Guided.
Base Prompt
This is the simplest prompt we employ to simulate a real-world scenario where an engineer has access only to a policy and requests. It consists of the policy that needs to be repaired in accordance with the must-allow and must-deny list specifications.
@@
We construct this prompt by first identifying the requests that each statement in a policy misclassifies. This augments the base prompt by providing the LLM with detailed information that maps specific failures to their root causes within the policy using an SMT-based fault localization report. Figure 3 illustrates an example. The prompt includes the policy, the must-allow and must-deny requests, and a fault-localization finding that indicates the error type, the failed requests, and the responsible statements for the failure.
@@
{Sid: VisualEditor0, Effect: Allow, Action: iam:GetAccountSummary, Resource: *}
{Sid: VisualEditor2, Effect: Allow, Action: s3:*, Resource: *}
{ Sid: VisualEditor3, Effect: Deny, Action: ec2:DescribeInstances, Resource: * }]}
@@
Resource: [arn:aws:s3:::admin-category/document.txt]
@@
Action: [sqs:SendMessage],
Resource: [arn:aws:athena:us-east-1:123456789012:workgroup/primary]
@@
Action: [sqs:SendMessage],
Resource: [arn:aws:s3:::admin-category/document.txt]
@@
Action: [ec2:DescribeInstances],
Resource:[arn:aws:glue:us-east-1:123456789012:table/my-database/my-table]
@@
Action: [ec2:DescribeInstances],
Resource:[arn:aws:athena:us-east-1:123456789012:workgroup/primary]
@@
Resource: [arn:aws:s3:::admin-category/document.txt]
@@
Responsible stmt: VisualEditor2
@@
Action: [sqs:SendMessage]
Resource: [arn:aws:athena:us-east-1:123456789012:workgroup/primary]
@@
SHOULD BE ALLOWED BUT EXPLICITLY DENIED
Action: [ec2:DescribeInstances]
Resource: [arn:aws:athena:us-east-1:123456789012:workgroup/primary]
Expected: allow, Got: deny
Responsible stmt: VisualEditor3
Figure 3: An example of a structured input provided to the LLM for a fault-localization-guided repair. The prompt is composed of three parts: the original faulty POLICY, the set of REQUESTS, and the FAULT LOCALIZATION report. The report explicitly maps failed requests to the policy statements responsible for the errors, categorizing them into three types of failures: "denied but explicitly allowed", "allowed but implicitly denied", or "allowed but explicitly denied".

### v2

@@ Abstract
Access control policies are vital for securing modern cloud computing, where organizations must manage access to sensitive data across thousands of users in distributed system settings. Cloud administrators typically write and update policies manually, which can be an error-prone and time-consuming process and can potentially lead to security vulnerabilities. Existing approaches based on symbolic analysis have demonstrated success in automated debugging and repairing access control policies; however, their generalizability is limited in the context of cloud-based access control. Conversely, Large Language Models (LLMs) have been utilized for automated program repair; however, their use for repairing such policies remains unexplored. In this work, we introduce CloudFix, the first automated policy repair framework for cloud access control that combines formal methods with LLMs. Given an access control policy and a specification of allowed and denied access requests, CloudFix employs formal methods-based Fault Localization to identify faulty statements in the policy and leverages LLMs to generate potential repairs, which are then formally verified using SMT solvers. To evaluate CloudFix, we curated a dataset of 282 real-world AWS access control policies extracted from forum posts and augmented them with synthetically generated request sets based on real scenarios. Our experimental results show that CloudFix improves repair accuracy over a baseline implementation across varying request sizes. Our work is the first to leverage LLMs for policy repair, showcasing the effectiveness of LLMs for access control and enabling efficient and automated repair of cloud access control policies. We make our tool and AWS dataset publicly available.
@@ I Introduction
Cloud compute systems allow administrators to write policies that specify access control rules, which govern how access to private data is granted. Policies are typically written in a convenient language, such as the Amazon Web Services (AWS) Identity and Access Management (IAM) policy language, allowing administrators to write their own policies. However, this is a tedious, complex, and often error-prone process. Moreover, policies must be modified on a regular basis due to the ever-changing nature of cloud computing. Creating and maintaining policies is a challenging endeavor, often requiring significant technical skills and domain expertise to ensure access control is correctly implemented. This can cause smaller organizations or those without the proper resources to forego best security practices in ensuring access privilege is kept to a minimum, and instead provide broad access to ensure their cloud computing services work as intended.
@@
Cloud computing services provide tools to assist administrators in writing and maintaining policies or monitoring access to resources (e.g., AWS IAM Policy Simulator, AWS CloudTrail service). However, when policies do not function as intended, or must be modified to incorporate access to new services or resources, these tools are inadequate when administrators lack the technical expertise. This is compounded by the lack of automated approaches for creating or repairing policies, as existing techniques either do not extend to cloud policies or have limited repair capabilities [25, 20, 4, 1, 13]. What is needed is automated techniques that can reason about the complexities of cloud policies and can modify and repair policies as permissions and services evolve throughout the cloud ecosystem.
@@
Recently, Large Language Models (LLMs) have seen great success in the field of software engineering [11]. They have become invaluable tools for software developers and engineers with their ability to generate code and automate tasks. In particular, LLMs are useful for automated program repair, where LLMs generate patches to fix faulty programs [30, 17, 10, 28, 3, 19]. In the field of access control, LLMs are successful in synthesizing policies from natural language descriptions [14, 26, 16, 24]. However, their applicability in the context of repairing access control policies has yet to be explored.
@@
In this work, we take the first step towards leveraging Large Language Models for access control policy repair. Given a faulty or misconfigured access control policy and access control requests that must be allowed or denied by the policy, our approach repairs the policy so that it correctly classifies the given set of requests. We implement our approach in a tool called CloudFix and showcase its effectiveness in repairing AWS IAM access control policies. Additionally, we curate the first policy repair dataset for cloud access control policies and make it publicly available for future research endeavors. Specifically, we make the following key contributions:
@@
  • We automate the access control policy repair process using an LLM-driven Fault Localization guided approach.
@@
  • We experimentally show that our Fault Localization techniques enable LLMs to repair faulty policies.
@@
The paper is organized as follows. In Section II we provide an overview of access control policies and the policy repair problem. In Section III we introduce CloudFix. In Section IV, we describe how we curate a dataset to evaluate CloudFix. In Section V we experimentally validate CloudFix. In Section VI, we discuss our findings. In Section VII we discuss the related work. In Section VIII, we conclude our paper.
@@ II Overview and Background
In this work, we focus primarily on access control policies written in AWS IAM policy language. However, the techniques we present in this work are not specific to the AWS IAM policy language and can be applied to any policy language.
@@ II-A Amazon Web Service Policies
An AWS IAM policy $P$ is a declarative policy consisting of a list of statements describing how access is governed. Policies field access control requests, which are either allowed or denied by the policy. An access control request is a tuple $(\delta,a,r,e)$ where $\delta$ is a specific principal performing the access, $a$ is the action to be performed, $r$ is the resource being accessed, and $e$ is the set of environment variables within the request (e.g., IP address, current time, etc). A policy statement $S$ is a 5-tuple (Effect, Principal, Action, Resource, Condition) where:
@@
Initially, all requests are implicitly denied. An access control request is granted access if and only if there is at least one statement within the policy that explicitly allows access and no statement in the policy explicitly revokes or denies access. The AWS policy language allows special characters to be used which represent broad access: the ’*’ wildcard character can be used to represent any string, while the ’?’ character can be used to represent any character. Additionally AWS allows identity-based policies, which omit the Principal field.
Figure 1: Manual Policy Repair Process. Stage 1: administrator authors policy, user generates requests. Stage 2: cloud engine validates requests, allowing s3:GetObject but implicitly denying ec2:RunInstances. Stage 3: The administrator adds an Allow statement and resubmits for validation.
@@ II-B A Motivating Example
In practice, debugging and repairing access control policies is largely manual. Administrators write policies in the AWS IAM console, then iteratively test them using the IAM Policy Simulator or CloudTrail logs until the desired permissions are achieved. A common workaround is to craft overly broad permissions (e.g., Action: s3:"*") to avoid unintended denials, then narrow scope iteratively—a risky approach.
Figure 1 illustrates this process. In Stage 1, an administrator authors a policy granting s3:GetObject access. A user then submits two requests: one for s3:GetObject (allowed) and one for ec2:RunInstances (implicitly denied, as no statement allows it). In Stage 2, the Cloud Engine validates these requests and returns the misclassification. In Stage 3, the administrator manually adds an Allow statement for the denied request and redeploys.
This process is tedious and error-prone. Each misclassification requires inspecting the policy line by line, tracing evaluation logic, and cross-checking CloudTrail logs. As the number of users, policies, and requests grows, manual repair becomes increasingly impractical.
@@ III CloudFix
In this section, we introduce our policy repair framework, CloudFix. It consists of four steps: (A) Goal Validation, (B) Fault Localization, (C) Prompt Generation, and (D) Repair Synthesis. Figure 2 shows the overall stages of the method. CloudFix operates over an input access control policy and request specification. The Goal Validator first evaluates the policy against the request specification: if all requests are correctly classified, the policy is considered correct, and the framework returns the policy. Otherwise, the process proceeds iteratively. In each iteration, the faulty policy and the requests are fed into the Fault Localizer component that analyzes the failed requests and the current policy to identify statements that are responsible for misclassifications. These faulty statements, together with the policy and requests, are provided to the Prompt Generator, which produces an input prompt for the Repair Synthesizer. The Repair Synthesizer queries an LLM to generate a candidate policy as a potential repair. This candidate policy is validated against the requests by the Goal Validator. If the candidate policy correctly classifies all requests, then the policy has been completely repaired, and the algorithm terminates. Otherwise, the candidate policy becomes the new input to the next iteration of the loop. This process continues until a fully repaired policy is produced, or the iteration threshold is reached. We retain the policy with the highest accuracy once the maximum iteration is reached. Alg. 1 presents the complete CloudFix algorithm.
@@ III-A Goal Validator
The Goal Validator determines if the policy correctly classifies all requests in the given specification. Alg. 2 shows how a request specification, $r$ in a set $R$ is validated against a policy $P$ to test whether $r$ is not allowed by $P$ and determine that $P$ is faulty. To do this, both the policy $P$ and $r$ are encoded into logical formulas suitable for satisfiability checking using the Quacky tool [8]. The algorithm returns FAIL if at least one request is misclassified. This information is then passed to the Fault Localizer for further identification of statements within $P$ that are responsible for the misclassification of the $r$ in $S$. Encoding the requests and the policies as SMT formulas using an SMT Solver enables us to eliminate the dependency on external policy evaluation engines. We use a maximum of $I$ iterations per policy to reach a repair goal of 100%. This can be defined by the user depending on the available resources. Once the target accuracy or the maximum $I$ is reached, the repair algorithm stops and returns the repaired policy. If the target repair accuracy is not achieved within $I$ iterations, the algorithm retains the repair with the highest accuracy.
@@
4:if $\mathit{verification}=\textsc{pass}$ then
@@
9:for each $(\delta,a,r,e)\in R_{\text{deny}}$ in $R$ do
@@ III-B Fault Localizer
Once the Goal Validator determines that a policy $\mathit{P}$ is faulty (i.e., it misclassifies requests), then $\mathit{P}$ is passed to the Fault Localizer to identify statements within it that misclassify the requests. We define a request misclassification as allowing a request that is supposed to be denied and/or denying a request that is supposed to be allowed by the policy. We define three types of faults within the policy that are responsible for misclassifications: case 1) faulty statements that explicitly allow requests which should be denied as per the request specification, case 2) statements that explicitly deny requests that should be allowed per the request specification, and case 3) requests which are implicitly denied yet should be allowed. The process is shown in Alg. 3. Each misclassified request is analyzed with respect to the policy to determine the type of fault that incorrectly classifies it.
Given a policy $\mathit{P}$ and a request set $\mathit{R}$, lines 5 - 19 in Alg. 3 illustrate how we identify faulty statements within $\mathit{P}$. The $\mathrm{SMTSolver}$ function determines the satisfiability of policy constraints against specific requests using the Z3 solver [6] for satisfiability checking and the Quacky tool for policy and request encoding. For each misclassified request, to check if it falls into case 1, $\mathrm{SMTSolver}(r,s)$ iteratively checks each allow statement $\mathit{s\in S^{+}}$ to see if it allows request $\mathit{r}$ by creating a minimal policy containing only that statement. Lines 8 and 9 show that if $\mathrm{SMTSolver}(r,s)$ returns $\mathit{SAT}$, we assert that there is an explicit allow statement that is allowing the statement that should have been denied. To check case 2, $\mathrm{SMTSolver}(r,P^{\star}\cup\{s\}))$ combines the universal policy with a deny statement, $\mathit{s}$, to determine if the deny statement explicitly denies $\mathit{r}$. This is shown on line 11. If the solver returns $\mathit{UNSAT}$, it indicates that the deny statement in $\mathit{S}$ is responsible for the incorrect denial of $\mathit{r}$. If this condition fails, case 3 holds, as shown in line 14. Finally, Alg. 3 returns a list that has a mapping between policy statements and request failures, along with the failure types. This combination of statements, along with the explicit requests they allow, deny, or implicitly deny, is then passed to the Prompt Generator, which is discussed below.
Figure 2: Architecture of CloudFix.
@@ III-C Prompt Generator
Once the Fault Localizer identifies the faulty statements $s$ within $P$, we format the results into a structured natural language prompt for the LLM. The Prompt Generator receives the Fault Localization report, the list of requests $R$, and the policy $P$. It then structures them into a format suitable for LLMs. We annotate the prompts with the iteration number and the accuracy obtained from the Goal Validator to provide contextual feedback to the LLM. In this work, we consider two types of prompts: baseline and Fault Localization guided.
Baseline prompt
This is a simple prompt we employ to simulate a real-world scenario where an engineer has access only to a policy and requests. It consists of the faulty policy that needs to be repaired in accordance with the must-allow and must-deny requests.
@@
We construct this prompt by first identifying the requests that each statement in a policy misclassifies. This augments the baseline prompt by providing the LLM with detailed information that maps specific failures to their root causes within the policy using an SMT-based Fault Localization report. Figure 3 illustrates an example. The prompt includes the policy, the must-allow and must-deny requests, and a Fault Localization report that indicates the error type, the failed requests, and the responsible statements for the failure. The complete prompts used in our experiments are provided in the accompanying artifact repository (see Section IX).
@@
{Sid: V1, Effect: Allow, Action: iam:GetAccountSummary, Resource: *}
{Sid: V2, Effect: Allow, Action: s3:Get*, Resource: *}
{Sid: V3, Effect: Deny, Action: s3:PutObject, Resource: * }]}
@@
Resource: [arn:aws:s3:::admin/doc1]
@@
Action: [s3:DeleteObject],
Resource: [arn:aws:s3:::user/data]
@@
Action: [s3:DeleteObject],
Resource: [arn:aws:s3:::admin/var]
@@
Action: [s3:ListBucket],
Resource:[arn:aws:s3:::admin]
@@
Action: [s3:PutObject],
Resource:[arn:aws:s3:::user/history]
@@
Resource:[arn:aws:s3:::admin/doc1]
@@
Responsible stmt: V2
SHOULD BE ALLOWED BUT EXPLICITLY DENIED
Action:[s3:PutObject]
Resource:[arn:aws:s3:::user/history]
Expected:allow, Got: deny
Responsible stmt: V3
@@
Action: [s3:DeleteObject]
Resource:[arn:aws:s3:::user/data1]
@@
Figure 3: Example of a Fault Localization guided prompt showing the faulty policy, requests, and fault localization report mapping failures to responsible statements.